# Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization

### v3

@@ IV-D Ablation Studies
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between drone and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with the baseline, the incorporation of DHML leads to significant performance improvements in both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in DVGL scenarios and enhances intra-view feature representation and discrimination.
TABLE V: The influence of each component on the performance of proposed method.
@@
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to intra-view information. To address this limitation, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%. These results validate the effectiveness of ICEL in enhancing cross-view consistency for DVGL.
Pseudo-Label Enhancement Refining Feature Learning with Robust Labels. To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability. After incorporating PLE, our final model achieves state-of-the-art performance, with R@1 increasing to 90.17% and 95.01%, and AP reaching 91.67% and 88.95%. This confirms that refining pseudo-label quality significantly enhances the final model’s effectiveness in cross-geographic view matching.
The ablation study systematically evaluates the contribution of each component in our method. Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline. On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement. Experimental results show that our method enhances the learning of discriminative and view-consistent representations under self-supervised learning.
Figure 4: The first row presents the t-SNE[54] visualization, while the second row illustrates the similarity distribution for randomly selected geographic locations. In the t-SNE visualization, each color represents a distinct location, where ’×’ markers correspond to the drone viewpoint, and the triangles indicate the satellite perspective of the same location.
IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the method, we perform a feature space and similarity distribution visualization for our method, as presented in Fig. 4. Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct that indicates better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our method.
Figure 5: Cluster analysis across different epochs for various datasets. The plot shows the number of clusters for DenseUAV, University-1652, and SUES-200 datasets.
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts in each epoch for three datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater drone image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (120 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates the ability of our method to learn discriminative representations progressively.
Figure 6: Top-6 retrieval results of the proposed method on the University-1652 dataset.
@@ IV-E Visualization
Compared with University-1652, the DenseUAV dataset introduces more realistic and complex challenges, such as dense spatial sampling, multi-scale resolutions, and temporal dynamics. In the drone $\rightarrow$ satellite retrieval setting, despite the high risk of confusion caused by dense spatial sampling, our method effectively captures spatial-geographic cues and achieves high retrieval accuracy, demonstrating robust cross-view performance under realistic and complex conditions.
Figure 7: Top-6 retrieval results of the proposed method on the DenseUAV dataset.
@@ V CONCLUSION
This paper investigates a highly valuable and challenging DVGL task to reduce reliance on paired drone–satellite images. To this end, we propose a self-supervised learning method that employs a shallow backbone with only 9 layers for feature extraction. Pseudo-labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module. Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across-views. To improve the quality and robustness of pseudo-labels, we introduce a pseudo-label enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, and even surpasses several fully supervised methods in certain scenarios without relying on paired drone–satellite images. These results validate the effectiveness and generalization capability of our method and verify the potential for real-world DVGL applications.

### v5

@@ IV-D Ablation Studies
TABLE IV: Comparisons between DMNIL and existing methods on DenseUAV
DenseUAV
All height 80m 90m 100m
Model Learning Type Parameters(M) R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
ConvNext-Tiny[DenseUAV] Supervised 30.10 60.23 81.94 - - - - - -
Sample4Geo[deuser2023sample4geo] Supervised 87.57 80.57 96.53 80.57 96.53 86.87 98.71 91.38 99.74
CAMP[wu2024camp] Supervised 91.40 88.72 97.76 76.45 96.01 84.81 99.23 89.19 99.23
DAC[xia2024enhancing] Supervised 96.50 84.47 96.53 80.92 97.04 85.32 98.71 85.46 99.10
MEAN[chen2024multi] Supervised 36.50 90.18 97.86 78.76 97.04 85.33 99.10 89.32 98.97
EM-CVGL[li2024learning] Self-Supervised 2.25 18.15 50.97 18.02 50.36 17.76 50.19 17.50 49.98
CDIKTNet[chen2025limited] Self-Supervised 91.52 22.57 40.24 17.25 40.15 23.04 48.39 25.48 52.38
DMNIL Self-Supervised 28.59 73.79 90.69 73.10 92.92 74.65 95.37 75.03 95.50
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the baseline, R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between drone and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with baseline, incorporation of DHML leads to performance improvements in both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48%, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances intra-view feature representation and discrimination.
TABLE V: Influence of each component on performance of DMNIL
@@
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, it still exhibits notable limitations in learning cross-view information. To address this, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 94.01%, while AP reaches 88.31% and 82.97%. These results validate the effectiveness of ICEL in enhancing cross-view consistency for DVGL.
Pseudo-Label Enhancement Refining Feature Learning with Robust Labels. To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability. After incorporating PLE, our final model achieves state-of-the-art performance, with R@1 increasing to 90.17% and 95.01%, and AP reaching 91.67% and 88.95%. This confirms that refining pseudo-label quality enhances the final model’s effectiveness in DVGL.
Figure 3: Accuracy of Pseudo-Labels. (a-b): Evolution of pseudo-label quality during training on the University-1652. (c-d): Evolution of pseudo-label quality during training on the DenseUAV datasets. “Pseudo-labels (w/o PLE)” and “Pseudo-labels (w/ PLE)” denote models trained without and with the proposed pseudo-label enhancement (PLE) strategy, respectively. The upper plots show the number of valid samples, correct and error pseudo-labels, and outliers obtained at each epoch, while the lower plots present the corresponding pseudo-label accuracy. The abbreviations “Max,” “Min,” and “Avg” in the table denote the maximum, minimum, and average values, respectively.
TABLE VI: Influence of DBSCAN EPS parameter applied to drone-view on University-1652.
@@
R@1 AP R@1 AP
Drone 0.1 384 35390 22.23 26.63 60.91 14.92
0.2 1249 26498 54.60 59.40 90.87 52.64
0.3 1785 15777 77.66 80.95 91.44 72.29
0.4 1656 7535 90.17 91.67 95.01 88.95
0.5 1100 2319 89.86 91.44 94.43 88.88
TABLE VII: Influence of DBSCAN EPS parameter applied to satellite-view on University-1652
[…]
R@1 AP R@1 AP
Satellite 0.1 831 8 89.29 90.89 94.57 87.74
0.2 827 8 89.77 91.32 94.43 88.78
0.3 827 8 90.17 91.67 95.01 88.95
0.4 821 8 88.72 90.41 94.43 87.28
0.5 811 0 89.64 91.18 94.72 88.10
TABLE VIII: Performance evaluation of different clustering algorithms on University-1652
[…]
Drone Satellite Drone Satellite R@1 AP R@1 AP
HDBSCAN[mcinnes2017hdbscan] 1260 818 5563 0 1.22 2.28 0.85 0.99
Agglomerative Clustering[ward1963hierarchical] 28697 840 0 0 29.54 29.54 67.18 29.59
DBSCAN[ester1996density] 1656 827 7535 8 90.17 91.67 95.01 88.95
IV-E Further Analysis
Clustering Analysis. We further conducted a sensitivity analysis of the DBSCAN parameters and performance comparisons of different clustering algorithms. As shown in Tables VI and VII, the parameter EPS influences both the number of clusters and outliers that directly impact model performance. For the drone-view, as EPS increases from 0.1 to 0.5, the number of clusters initially rises and then falls, while outliers decrease sharply. The best performance is achieved at EPS = 0.4. When EPS becomes too large, dissimilar samples are incorrectly merged, which leads to a slight performance drop. For the satellite-view, results remain relatively stable, with the best performance at EPS = 0.3. As shown in Table VIII, among three different clustering algorithms, HDBSCAN requires no manual tuning but yields many small clusters and lower accuracy. Agglomerative Clustering groups nearly all samples but performs suboptimally. In contrast, DBSCAN filters noisy samples and form compact, discriminative clusters, which achieves the best performance.
Analysis of the Accuracy of Pseudo-Labels During the Training Stage. We further investigate the variations of pseudo-labels during training on the University-1652 and DenseUAV datasets, as shown in Fig. 3. On University-1652, the number of valid samples rapidly increases and stabilizes at about $33\times 10^{3}$. The pseudo-label accuracy reaches 95.63% at its peak and maintains an average of 88.46%. With the proposed PLE module, both the valid samples and accuracy are further improved, while the number of erroneous pseudo-labels gradually decreases during training. On DenseUAV, a similar trend can be observed. Without PLE, valid pseudo-labels stabilize around $22\times 10^{3}$ and the accuracy reaches 89.20% with an average of 82.47%. After applying PLE, valid samples increase to $23\times 10^{3}$ and the accuracy improves to 92.83% with an average of 88.31%. These results confirm that PLE effectively improves pseudo-label quality.
Analysis of Perturbation Effect. To further evaluate the influence of perturbation in the PLE, we compare the performance under settings with and without perturbation. As shown in Table IX, the model trained without perturbation achieves stable but slightly lower retrieval accuracy. Introducing moderate Gaussian perturbation ($\sigma=0.05\sim 0.15$) consistently improves the results in both retrieval directions, indicating that the perturbation facilitates smoother feature alignment and enhances local consistency during training. The best performance is achieved at $\sigma=0.10$, which provides an optimal trade-off between feature diversity and stability. In contrast, an excessive perturbation level ($\sigma=0.20$) introduces noise that disturbs the learned representations, leading to performance degradation.
TABLE IX: Performance comparison with and without perturbation.
[…]
R@1 AP R@1 AP
w/o Perturbation - 89.49 91.10 94.15 87.41
w/ Perturbation 0.05 90.17 91.67 95.01 88.95
0.10 90.45 91.94 95.15 88.85
0.15 90.37 91.92 95.15 88.74
0.20 89.20 90.89 94.02 86.00
Hyper-parameter Analysis for $\alpha$ and $\xi$. The influence of hyper-parameters $\alpha$ in Eq.(4) and $\xi$ in Eq.(8) are shown in Fig. 4. In the baseline, which learns from cluster-level representations, $\alpha$ is set to 0.1, as adopted in previous works [dai2022cluster, adca]. When the momentum factor increases in Eq. (4), the update of the current query instance feature becomes slower, which weakens its adaptability to new samples and feature variations. In the DHML module, historical consistency is emphasized by maintaining a 7:3 ratio between historical and current features. This design balances stability and adaptability, which leads to the best performance when $\xi=0.3$. This discrepancy arises from the fundamentally different motivations of momentum factor in the two designs. The baseline requires fast and responsive feature updates to track the evolving cluster assignments. In contrast, DHML aims to accumulate stable historical information and suppress noisy in early stage.
Hyper-parameter Analysis for $k_{1}$ and $k_{2}$. As shown in Fig. 5, the parameter $k_{1}$ controls the size of a strictly filtered, high-confidence neighborhood, while $k_{2}$ defines an expanded neighborhood that captures additional informative samples ($k_{2}>k_{1}$). When $k_{1}$ is too small, the neighborhood lacks sufficient diversity and informative positives, whereas a large $k_{1}$ introduces low-confidence samples and noise. Similarly, increasing $k_{2}$ enhances cross-view contextual relations, but excessive expansion may include irrelevant neighbors and lead to slight degradation. In this work, $k_{1}$ and $k_{2}$ are set to 10 and 20, respectively. It is noted that slightly adjusting these parameters may lead to further performance improvements.
Figure 4: Performance influence of momentum factors. (a) Performance influence of the momentum factor $\alpha$ in the baseline on the University-1652. (b) Performance influence of the momentum factor $\xi$ in the DHML module on the University-1652.
Figure 5: Performance influence of the parameters $k_{1}$ and $k_{2}$ .(a) Performance influence of the $k_{1}$ in the ICEL module on the University-1652 dataset. (b) Performance influence of the parameter $k_{2}$ in the ICEL module on the University-1652 dataset.
Analysis of Robustness. To further evaluate the robustness of DMNIL, the analysis considered sparsity with respect to both features and samples. For feature sparsity, we randomly masked three black patches on each test image to simulate missing visual information and reduce cross-view overlap. The patches were allowed to overlap. The ratio of the masked area to the total area was controlled by a parameter $A_{s}$. For sample sparsity, since each location in the original test set contains only one satellite image but multiple drone images, we controlled the number of available drone images per location to simulate varying degrees of sample sparsity. The corresponding results are reported in Table  X and Table  XI.
For feature sparsity, as shown in Table X, when increasing the occluded area ratio $A_{s}$ from 0.1 to 0.5, both R@1 and AP of DMNIL decrease gradually but remain consistently higher than those of the baseline. Even at $A_{s}=0.5$, where half of the visual content is randomly masked, DMNIL still achieves 36.88% R@1 and 41.63% AP, outperforming the baseline by a considerable margin. For sample sparsity, as shown in Table XI, the performance of both methods declines as the number of available drone images per location decreases from 20 to 1. Nevertheless, DMNIL consistently achieves higher retrieval accuracy than the baseline, reaching 91.16% R@1 with only a single drone image. The relatively small performance gap under reduced samples verifies that DMNIL generalizes well even when the cross-view matching pairs are extremely limited.
[…]
As (%) DMNIL Baseline
R@1 AP R@1 AP
0.1 84.69-5.480 86.99-4.680 64.67-6.090 69.13-5.780
0.2 72.89-17.28 76.34-15.33 52.54-18.52 57.70-17.21
0.3 58.89-31.28 63.11-28.56 40.12-30.94 45.22-29.69
0.4 45.27-44.90 49.98-41.69 30.22-40.84 35.07-39.84
0.5 36.88-54.29 41.63-50.04 23.76-47.30 28.31-46.60
[…]
Number of Drone DMNIL Baseline
R@1 AP R@1 AP
1 91.16-3.990 92.44+3.590 66.61-14.70 70.52+5.260
5 93.56-1.590 89.07+0.220 76.03-5.280 66.68+1.420
10 93.72-1.430 89.11+0.260 77.74-3.570 65.60+0.340
15 94.29-0.860 89.15+0.300 78.60-2.710 65.58+0.320
20 95.10-0.050 89.20+0.350 79.60-1.710 65.43+0.170
TABLE XII: Different configurations of learning type on University-1652
University-1652
[…]
Model Learning Type Parameters(M) R@1 AP R@1 AP
Backbone Supervised 28.59 90.12 91.96 94.92 89.72
DMNIL Self-Supervised 28.59 90.17 91.67 95.01 88.95
Supervised 28.59 92.02 93.36 95.57 91.27
TABLE XIII: Different configurations of learning type on SUES-200
SUES-200
$\rightarrowSatellite Drone\rightarrowSatellite$
150m 200m 250m 300m 150m 200m 250m 300m
Model Learning Type Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP R@1 AP
Backbone Supervised 28.59 87.80 87.80 94.00 95.22 96.52 97.19 96.32 96.32 96.25 87.22 97.50 92.94 98.75 94.92 97.50 94.66
DMNIL Self-Supervised 28.59 87.00 89.20 93.98 95.11 96.50 97.23 97.23 97.86 95.00 85.46 97.50 92.72 100.00 95.43 100.00 96.39
Supervised 28.59 94.00 95.21 97.83 98.88 99.14 95.44 99.25 99.39 97.50 92.71 97.50 97.43 100.00 98.98 98.75 98.25
[…]
Figure 7: Feature Space and Similarity Distribution. The first row presents the t-SNE[van2008visualizing] visualization, while the second row illustrates the similarity distribution for randomly selected geographic locations. In the t-SNE visualization, each color represents a distinct location, where ’×’ markers correspond to the drone view, and the triangles indicate the satellite-view of the same location.
Analysis of Unsupervised Learning. To further explore the advantages of self-supervised learning, three configurations are reported for comparisons. In supervised learning type, the backbone of DMNIL is trained solely with an InfoNCE loss using paired data. In self-supervised learning type, the DMNIL employs the same backbone that is trained by $\mathcal{L}_{total}$ in Eq. (23) without any paired labels, which is namely the proposed method in this work. Furthermore, in another supervised learning type that inspired by [38], the DMNIL employs the same backbone trained with InfoNCE and $\mathcal{L}_{total}$ using both paired data and unpaired data. As shown in Tables XII and XIII, the supervised learning type of backbone achieves performance comparable to the proposed method in this work, which benefits from the paired relationship that is unavailable in self-supervised scenarios. It is crucial that the proposed method attains this accuracy without requiring pair data, which has demonstrated that its capacity to effectively exploit latent information from unpaired data. It is a significant advance for real-world applications where paired data is severely scarce.
The supervised type of DMNIL further enhances the performance and obtains the latent information outside of the paired relationship. This demonstrates that the information learned from unpaired data is not equivalent to the paired data, and their combination gives additional performance improvement. Nevertheless, the fact that self-supervision can extract latent information from unpaired data highlights its unique value, in which such information cannot be fully replaced by explicit pairing cues. This phenomenon is highly interesting, and we will further explore the mechanisms behind it in future works.
IV-F Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the method, we perform a feature space and similarity distribution visualization for DMNIL, as presented in Fig. 7. Comparing (a) Baseline and (b) DMNIL, it is evident that DMNIL enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct, which indicates better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. DMNIL demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of DMNIL.
[…]
@@ V CONCLUSION
This paper investigates a highly valuable and challenging DVGL task to reduce reliance on paired drone–satellite images. To this end, we propose a self-supervised DMNIL method that employs a shallow backbone with only 9 layers for feature extraction. Pseudo-labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. We design a DHML module to further improve intra-view discriminability and consistency. Additionally, to mitigate the structural and semantic domain gap between views, we propose an ICEL module, which facilitates semantic alignment by learning high-dimensional feature consistency across-views. To improve the quality and robustness of pseudo-labels, we introduce a pseudo-label enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed DMNIL method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks, including University-1652, SUES-200, and DenseUAV, demonstrate that DMNIL outperforms existing self-supervised methods, and even surpasses several fully supervised methods in certain scenarios without relying on paired drone–satellite images. These results validate the effectiveness and generalization capability of DMNIL and verify the potential for real-world DVGL applications.